OpenStudy (anonymous):

Using the Intermediate Value Theorem, show that the function f has a zero between a and b. f(x)=x^3+4x^2-8x-10 a=-6, b=-5 What is f(-6)? What is f(-5)? Is there a zero between x= -6 and x= -5?

OpenStudy (anonymous):

f(-6) means you substitute in -6 everywhere you see an x in the function. f(-5) means you substitute in -5 everywhere you see an x in the function. Once you compute those values, if one is positive and one is negative, what must happen to the function between those two numbers?

OpenStudy (anonymous):

i have no idea :S

OpenStudy (anonymous):

Okay, I'll show you -6. f(-6) = (-6)^3 + 4(-6)^2-8(-6)-10 = -216 + 144 + 48 - 10 = -34

OpenStudy (anonymous):

how are you getting 4(-6)^2 = 144? sorry im really bad at this

OpenStudy (anonymous):

(-6)^2 = (-6) * (-6) = 36 Multiply 36 by 4 = 144